OpenStudy (anonymous):

Solve the quadratic equation. 3x^2 = 27 A. x = 3 and x = 27 B. x = 1 and x = 9 C. x = ±3 D. x = ±√3

OpenStudy (ucreategames):

c

OpenStudy (ucreategames):

27 /3=9. the sqrt of 9 is -3 and 3

mathslover (mathslover):

Hello Awesome_Naya , To find x, take square root both sides : \(\sqrt{3x^2} = \pm \sqrt{27}\) \(x\sqrt{3}= \pm \sqrt{9 * 3}\) \(x\sqrt{3} = \pm 3\sqrt{3}\) \(\implies x =\pm 3\)

OpenStudy (anonymous):

thank you